Home‎ > ‎Reports‎ > ‎Crew102‎ > ‎

day07


Crew 102 Log Book for April 1, 2011


Photos of the day





Commander's Check-In Report

Franco Carbognani Reporting

Crew Physical Status: good except for one which cold is slowly improving

Time Departed/Returned from EVA: 12:10 - 14:00, 17:00 - 20:00

Brief Narrative of Field Mission Results:

Performed tests on soil permeability

Discovered and GPS traced sites abundant in cyanobacteria samples

EVA Data/Interpretations:

Rock Samples analysis ongoing

EVA GPS tracing using crew smart-phones is working well

Engineering/Hab Maintenance:

Main generator running fine after provisional fixes

Water Heater leaking problem manageable keeping heating period to the minimum.

Plans for Tomorrow:

Long EVA with ATV for localisation of Gryphea Arcuata bio erosion samples sites

Short EVA around the Hab dedicated to corrosion inspection and sample collection for optical microscopy analysis

Proceed on analysis of Radiotelescope data after calibration. Verify the

possibility to define online analysis able to detect energetic solar storms.

Possibly proceeding on bringing back gray water system in operation based on the Mission Support Chief Engineer inputs.

Continue on ongoing psychological and physiological studies.

Support Requested:

Don Lusco for final fixes on power generator.


Commander's Report

Franco Carbognani Reporting

The provisional fix to the power outage problem is holding.

Don did not show up also today for a final fix but it seems the main generator is running fine for the moment.

A short EVA has been performed in the Hab vicinity for collecting geological sampling data.

A third long ATV EVA has been performed and the Crew is well advancing within the goal of tracing, via GPS coordinates, all the "sweet" cyanobacteria spots around for our and for the next MDRS crews studies.

Radiotelescope data collection for off-line analysis is proceeding smoothly.

Hub Monitoring and Alarming System up and running in provisional mode with corresponding web pages available for internal access. Waiting for feedback on how to proceed.

The straining tank in the Green Hab should be now functional. Waiting for instruction on the possibility to bring back the whole gray water pipeline

The Telescope mount and controller are ready in the Observatory to be handed over to Don Lusco when he will be on-site.

Mood of the crew is good.

John cold is slowly getting better but his cough is still very annoying and prevent him to sleep well.

It seems the season had suddenly switched from Winter to Summer. No need anymore today for the heating.

During today EVA near the Hab temperatures probed were over 30 C°

On water issue side, the explanation, proposed by Jean Hunter, that the white flakes are a manifestation of "temporary hardness" makes plenty of sense to us.


Engineering Report

Franco Carbognani Reporting

Generator/Electricity:

Kitty (Diesel Generator) (Used? Yes/No): Yes

Honey (Gas Generator) (Used? Yes/No): No

Battery Charge Level:

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

Kitty provisional repair seems holding. The cooling fan is running smoothly as the whole Kitty engine.


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0.65

Propane Tank (Full/0.75/0.5/0.25/Empty): 53%

Gasoline Tank (Full/0.75/0.5/0.25/Empty): 0.65

Notes/Comments:


Oil:

Kitty Oil Quantity (# of Quarts in storage): 4/4

Honey Oil Quantity (# of Quarts in storage): 0/4, to be refilled

ATV Oil Quantity (# of Quarts in storage): all 4/4

Car Oil Quantity (# of Quarts in storage): 3/4


Water Status:

Outside Potable Water Tank Level (inches from bottom): 18

Trailer Potable Water Tank Level (inches from bottom): 18

Hab Potable Water Tank Level (inches from bottom): 45

Potable Water Meter Reading: 44329,7

In to GreenHab Meter Reading:

Notes/Comments:

Water heater leaking problem: By greatly limiting the heating time we have seen no recurrence of the problem in the last 2 days

The Trailer Potable Water Tank is now almost less than half filled. May be worth planning for refilling


GreenHab:

Condition of plants in Tank 1:

Condition of plants in Tank 2:

Condition of plants in Tank 3:

Greenhab Notes/Comments:

Following Joshua detailed instruction we should have brought back the straining tank into operation.

Waiting for further instructions on how to proceed from there.


Observatory:

Telescope Used? (Yes/No):No

Observatory Notes/Comments:

Boxes containing the telescope mount and controller are in the Observatory in order to be handed over to Don Lusco


Transportation:

Hab Car (NOT RENTAL)

Used (yes/no): no

Oil Checked (yes/no):yes

Tire Status: good

Odometer at end of day: 123761.4

Notes/Comments on Hab Car:


ATV 1 (Opportunity):

Used: yes

Oil Checked: n

Fuel Consumed: 1/2 liter

Tire Status: inflated

Comments:


ATV 2 (Spirit):

Used: no

Oil Checked: n

Fuel Consumed: none

Tire Status: inflated

Comments: this ATV does not start


ATV 3 (Viking I):

Used: yes

Oil Checked: n

Fuel Consumed: 1/2 liter

Tire Status: good

Comments:


ATV 4 (Viking II):

Used: yes

Oil Checked: n

Fuel Consumed: 1/2 liter

Tire Status: good

Comments:


Heating and Ventilation:

Thermostat setting upstairs (Farenheit): 65

Actual temperature upstairs (Farenheit): 79

Thermostat setting downstairs (Farenheit):55

Actual temperature downstairs (Farenheit): 75


Computers/Networking Infrastructure:

Notes/comments on internet/computers?:


General Engineering Notes/Comments:



Science Report (Biology)

Lara Vimercati Reporting
 

HABITAT WATER RECYCLING SYSTEM TO GROW EDIBLE CYANOBACTERIA AT MARS DESERT RESEARCH STATION (MDRS)

The development of space programs that aim for long-term missions are recently arising the question on how humans are going to sustain their nutritional needs. Healthy food that meets all human nutrition requirements as well as a simple and inexpensive method to produce it in space will be key elements in determining the best food source. Due to its high nutrient content and its basic growth needs (light as energy source and CO2 as carbon source), Spirulina (Arthrospira platensis) is currently considered one of the best candidates for in situ food production for future long-term space missions.

Spirulina is a cyanobacterium that has long been known for its properties as a nutritious food source and health aid. Nearly 62% of its mass is made of proteins and it produces all the essential aminoacids for life. It contains  a wide range of micronutrients and also supplies high levels of antioxidants. Furthermore it is endowed with antiviral and antibacterial properties that might boost immune function by high level of antibodies production.

NASA found it to be  an exceptional compact food source (1 kg of Spirulina is equivalent to a 1000 kg of assorted vegetables) and is currently trying to implement methods to grow it for future missions.

Spirulina is known to need a high degree of water quality and this last one is going to be a scarce resource during long-term flights. Hence methods need to be developed to treat human wastewater in an efficient manner. An optimal method to recycle water will ultimately lay the foundations for mass production of Spirulina inside space habitats . Being a photosynthetic organism, Spirulina might also eventually address the problem of air recycling by providing oxygen and removing CO2 produced by astronauts. An enclosed fully functional enclosed ecosystem could then be proposed.

Such a self sustaining ecosystem would improve long-term space missions and this option could be fundamental for a successful trip to Mars.

A simulation of a manned mission to Mars with 6 crew members is currently being carried out at the Mars Desert Research Station (MDRS) in Utah.  An experimental program will be set up to evaluate the ability of forward osmosis to purify wastewater produced during a manned mission to Mars and to use it to grow Spirulina cultures. This study builds upon the current Habitat Water Wall for water and atmosphere recycle project being developed by NASA Ames Research Center. This project envisions a system of membranes embedded into the walls of an inflatable habitat structure used to recycle water to be reused by humans.

The experiment will be set up as follows: urine samples collected from the 6 crew members and greywater samples will be filtered using the commercially available forward osmosis technology X-Pack  (Hydration Technologies). Approximately one liter of each sample will be processed and the filtered water will be used to make growth medium for Spirulina (Schlosser, 1982).  Water quality will be tested measuring the chemical oxygen demand (COD) . A control culture of Spirulina will be grown in parallel using sterile DI water to make growth medium. All the cultures will be grown for 10 days inside the Green Hab of the MDRS station under diffused illumination at room temperature (28°C-32°C).Test cultures and control will be monitored daily and growth will be measured as a function of turbidity. pH will be monitored as well and stabilized between 8-11 units.

By the end of the experiment cells will be harvested and wet mass production will be compared between test cultures and control. Furthermore,  chemical analyses will be performed on the filtered samples.


EVA 6 Report

Kavya k.Manyapu Reporting

Time of Departure: 12.10 pm
Time of Arrival: 1.57 pm
Crew Members: K.Manyapu, F.Carbognani
Site Location:
Latitude: 38ᵒ24’23”
Longitude: -110ᵒ47’30”
Altitude: 4500 feet

Total Distance: 0.5 miles
Transit Mode: By foot
Objectives: The purpose of this EVA was to perform soil sampling and to conduct experiments to measure the soil permeability and soil porosity for the various types of soils in this area. The experiment also involved the measurement of temperature and humidity of the soil at various depths ranging from the surface to 1 foot deep.

Results: The soil samples of two different soil types were collected. The experiment to measure the soil permeability was conducted in two different areas with two different types of soil. The measurements were performed at 10 minute intervals lasting up to an hour. The crew was also successful in obtaining the temperature and humidity levels of the soil. The soil samples collected will be used to measure the porosity and the grain size in the habitat laboratory.

Lessons Learned:
Find a smarter way to measure humidity at depths more than 1 foot below the soil surface without disturbing the soil (In our case we had to dig holes of 1 foot).
This experiment is very time consuming, however more samples and measurements for soil permeability over a longer period of time might produce an accurate trend in the variation of temperatures, humidity as well as soil permeability (soil permeability measurements to be done in various areas around habitat because of the high variability in the type of soils in this area).
The spacesuit air management system and the bubble hats need to be replaces as soon as possible for greater visibility as well as air management since the crew finds it very hard to breath. With the strenuous activities the crew performs, optimal oxygen supply is required for the crew to be able to breathe well and have maximum productivity during EVAs.


EVA 7 Report

Yuval Brodsky Reporting

Time: 1700 - 2000

Crew Members: Y. Brodsky (EVA CDR), A. Kapoglou, L. Vimercati

Site Location: 12S 0518232 4255758

Transit Mode: ATV

Objectives: Cyanobacteria sample collection, site mapping

Results: ~400 grams of samples containing cyanobacteria were collected, with locations mapped.

Lessons Learned: N/A


Chef's Report

Dr. John Deaton Reporting

Time:  22:55

1a) Today was a cooking day

1b) Was today a special day celebrated at the main meal with special food or activity?

Last evening (March 31) crew celebrated watching a movie and consuming popcorn.  Was too late to put into March 31 report, but worthy enough to include here.

2)  List the foods served at today's main meal (usually the evening meal), giving full names of each.

Shepherd’s Pie (ground beef and mashed potatoes)

Stir fried tofu

Mixed vegetables

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)

Left over ground beef will be used tomorrow (perhaps breakfast burritos)?

4)  Recipes for anything more complicated than following package directions exactly, or rehydrating.

Nothing special; followed directions contained in menu book in the MDRS and past experience

 5)  Lessons learned

Corn is better if you let it soak before cooking

 6) Comments/questions for Kim and Jean, the food study investigators

Crew seemed to enjoy this meal.


Journalist Report

Kavya Manyapu Reporting

We have some of the smartest and brave engineers and scientists on board Crew 102 who have successfully completed a week on the Red Planet . Congratulations Crew 102!

The first EVA for the day was commanded by K.Manyapu accompanied by F.Carbognani. L.Vimericati served as the cap-com for the EVA. The objective of the EVA was to measure the soil permeability, porosity , temperature and humidity. The EVA lasted 1 hour and 43 minutes.  The second EVA for the day was a continuation of the cyanobacteria sample collection from the previous days EVAs and marking the areas where they are available on Mars (Analogue). This EVA was commanded by Y.Brodsky, accompanied by Crew Biologist L.Vimericati as well as A.Kapoglou which lasted 2 hours. The cap-comp for this EVA was F.Carbognani.  Some crew members received training and guidance in using the ETVs by Y.Brodsky. The rest of the day was spent analyzing data for the science projects. Fortunately, today was a cooking day (try to cook the dehydrated food using your own culinary creativity) and the crew seems to enjoy the cooking days more than the non-cooking days (pre-prepared food). Overall, it has been a tiring day for the crew with strenuous EVAs and experiments.

Hope the crew gets enough rest tonight to be fit and ready for tomorrow’s adventure (The Red Planet seems to have many surprises in store for the crew).

Ciao for now from Crew 102.